OpenStudy (anonymous):

2x – 3y + z = –1 2x – 2y – 2z = –8 x + 5y – 3z = 2 A.(–1, 2, 8) B. (1, 5, –3) C.(1, 1, 0) D.(1, 2, 3)

jhonyy9 (jhonyy9):

2x-3y+z=-1 2x-2y-2z=-8 ------------- 0 -y+3z=7 so than y=3z-7 2x-2y-2z=-8 x+5y-3z=2 multiply by -2 and assume 2x-2y-2z=-8 -2x-10y+6z=-4 ---------------- 0 -12y+4z=-12 than -12(3z-7)+4z=-12 -36z+84+4z=-12 -32z=-96 z=-96/(-32) = 3 z=3 so than y=3*3-7=9-7=2 y=2 2x-2y-2z=-8 2x-2*2-2*3=-8 2x=-8+4+6 2x=2 x=1 (1,2,3) will be right sure

jhonyy9 (jhonyy9):

D.

jhonyy9 (jhonyy9):

hope that is understandably su

jhonyy9 (jhonyy9):

sure

jhonyy9 (jhonyy9):

give me medal please thank you

jhonyy9 (jhonyy9):

good luck bye